Consider the following problem faced by a Dietician: A diet needs to be created that contains not less than 2213 calories, not more than 75 grams of protein, not less than 16 grams of carbohydrates and not less than 13 grams of fat. Also, the diet should have minimal cost. In addition the diet should include at least 1.1 Units of fish and at least 0.7 cup of milk. The diet will consist of the six different foods: Bread, Milk, Cheese, Fish, Potato and Yogurt. The following table lists the cost per unit of each item and the number of grams of protein, fat, and carbohydrates per unit as well as the calories that will be supplied by each food unit. Milk 23 9 7 16.7 178 Fish 81 8 8 0 205 Bread Cheese P Cost, (cents)8 6 3 15 25 7 15 0.4 182 Potato 15 1.3 0.1 30.6 114 Yogurt 29 11.2 3 25 342 Protein, g Fat, g Carbohydrates,g Calories, Ca176 Find a diet that meets the requirements above while keeping the cost at an absolute minimum. Make sure that your Food Units are all Integers
Answer #1

Let’s create a linear programming model and solve it on excel solver.

Here our decisions are the amount of bread, milk, cheese, potato, fish and yogurt to consume. We can denote them as X1, X2, X3, X4, X5, and X6.

The objective is to minimize the cost. Thus our objective function is

Minimize 8X1 + 23X2 + 25X3 + 15X4 + 81X5 + 29X6

The constraints are

176X1 + 178X2 + 182X3 + 114X4 + 205X5 + 342X6 >= 2213 (Calorie constraint)

6X1 + 9X2 + 7X3 + 1.3X4 + 8X5 + 11.2X6 <= 75 (protein constraint)

15X1 + 16.7X2 + 0.4X3 + 30.6X4 + 0X5 + 25X6 >= 16 (carbohydrate constraint)

3X1 + 7X2 + 15X3 + 0.1X4 + 8X5 + 3X6 >= 13 (fat constraint)

X5 >= 1.1 (fish constraint)

X2 >= 0.7 (milk constraint)

X1, X2, X3, X4, X5, X6 >= (non-negative constraint)

The solver model is shown below

-SUM(B18:G18) Bread Milk Cheese Potato Fish Yogurt 29 11.2 25 176 178 182 114 205 342 X6 23 25 15 81 2 Cost 3 Protein 4 Fat 5 Carb 6 Calories 15 15 16.70.4 30.6 X1 X2 X4 X5 0 9 Dec variables 0 10 11 Calorie const 0 12 Protein const 0 13 Carb const 14 Fat const 15 Fish const 16 Milk const Limit 0 2213 0 <= 75 0 16 013 01.1 0 0.7 Total 0 0 0 18 Cost 19 21 23 24 Sheet450 Sheet451 Sheet452 Sheet453 Shect454 + 11596 囲回凹

The formulas are shown below

The solver parameter is shown below

SUM(B18:G18) Solver Parameters G HK Bread Milk Cheese Potato Fish Yogurt 2 Cost 3 Protein 4 Fat 5 Carb 6 Calories 23 25 15 1.

The result is shown below. The total cost is 292 cents

Each items are

Bread = 6

Milk = 1

Cheese = 0

Potato = 2

Fish = 2

Yogurt = 1